Erile:

a 1280 lb white horse is cantering at 20 ft/s. Its linear momentum is

Erile:

What is the answer?

lowkey:

Physics is pointless, don't worry

Erile:

How to solve that i need a concrete answer

Shadow:

\[p = mv\] Momentum = mass times velocity \[p = 1280lb(20ft/s) = \frac{ 1280lb \times 20ft }{ s }\]

Shadow:

They may want you to convert ft to meters, in such case.. \[20ft \times \frac{ 0.3048m }{ 1ft }\]